### v1

# Preserving Expert-Level Privacy in Offline Reinforcement Learning

Abstract
The offline reinforcement learning (RL) problem aims to learn an optimal policy from historical data collected by one or more behavioural policies (experts) by interacting with an environment. However, the individual experts may be privacy-sensitive in that the learnt policy may retain information about their precise choices. In some domains like personalized retrieval, advertising and healthcare, the expert choices are considered sensitive data. To provably protect the privacy of such experts, we propose a novel consensus-based expert-level differentially private offline RL training approach compatible with any existing offline RL algorithm. We prove rigorous differential privacy guarantees, while maintaining strong empirical performance. Unlike existing work in differentially private RL, we supplement the theory with proof-of-concept experiments on classic RL environments featuring large continuous state spaces, demonstrating substantial improvements over a natural baseline across multiple tasks.
1 INTRODUCTION
Figure 1: Training pipeline for Expert-level Differentially Private Offline RL given an expert set $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{m}\}$ and input trajectories logged into an offline dataset $D_{\Pi}$. $\mathcal{A}_{DR}$ (Algorithm 2) splits input trajectories and adds stable prefixes to $D^{\Pi}_{stable}$ discarding the rest to $D^{\Pi}_{unst}$. These are used to train any off-the-shelf offline RL algorithm as described in Algorithm 3 and learn an expert-level private policy $\pi_{private}$.
Leveraging existing offline datasets to learn high-quality decision policies via offline Reinforcement Learning (RL) is a critical requirement in many settings ranging from robotics (Fu et al.,, 2020; Levine et al.,, 2020) and recommendation systems (Bottou et al.,, 2013; Ie et al.,, 2019; Cai et al.,, 2017) to healthcare applications (Oberst and Sontag,, 2019; Tang et al.,, 2022). Correspondingly, there is now a rich literature on effective techniques (Fujimoto et al.,, 2019; Kumar et al.,, 2019, 2020; Cheng et al.,, 2022) for learning from such offline datasets, collected using one or more behaviour policies. An often overlooked aspect of this literature is, however, that in privacy sensitive domains such as personalized retrieval, advertising, and healthcare, the behaviour policies might reveal private information about the preferences / strategies used by the corresponding experts (users, advertisers, health care providers) whose decisions underlie the offline data. In such scenarios, we ideally seek to uncover the broadly beneficial strategies employed in making decisions by a large cohort of experts whose demonstrations the offline data is collected with, while not leaking the private information of any given expert. We call this novel setting offline RL with expert-level privacy, and design algorithms with strong privacy guarantees for it.
There has been a growing interest in differentially private RL, both in online (Wang and Hegde,, 2019; Qiao and Wang,, 2023; Zhou,, 2022; Liao et al.,, 2023) and offline (Qiao and Wang,, 2024) settings. The offline RL literature, which is most relevant to this work, primarily focuses on protecting privacy at the trajectory-level. However, protecting the privacy of an expert, who might contribute multiple trajectories to the demonstration dataset, requires significantly more careful techniques. Secondly, the prior works mostly focus on tabular or linear settings (Qiao and Wang,, 2023, 2024; Zhou,, 2022), where the offline RL problem can be effectively solved using count or linear regression based techniques, and is hence amenable to existing techniques for privatizing counts or linear regression models. In contrast, we make no assumptions on the size of the state space, the parameterization of the learned policy or the policy underlying expert demonstrations.
Contributions: With this context, our paper makes the following contributions:
  • We identify and formalize the question of offline RL with expert-level privacy and motivate it with practical examples. We also explain the inadequacy of prior algorithms in this setting.
  • We provide practical algorithms for offline RL with strong expert-level privacy guarantees, building on top of any gradient-based offline RL algorithm. In contrast with prior approaches which use noisy statistics of the data in offline RL (Qiao and Wang,, 2024), we develop a novel scheme which also identifies a subset of the data which can be used in learning without any added noise. This capability, which we describe in more detail below, is not feasible in the typical trajectory-level privacy settings, and highlights the benefits of systematically studying the expert-level privacy setting.
  • We perform a proof-of-concept evaluation of our algorithm on standard RL benchmarks and compare our approach with an adaptation of DP-SGD to expert-level privacy as a baseline for privatizing the gradient-based offline RL algorithm. We show that our approach significantly improves upon this baseline across multiple test environments. The full results are discussed in Section 6.
Our algorithm: Our algorithm proceeds in two stages. The first is a data filtering stage that produces a set of trajectory prefixes, which can be used in training without any further noise addition, in a privacy-preserving manner. The intuition for this set is that it is quite likely to occur across many experts, and hence using it in training does not violate the privacy of an individual expert. The second part of the algorithm is to run DP-Stochastic Gradient Descent (DP-SGD) [^1] DP-SGD is to an SGD based approach where gradients coming from individual experts are clipped, and appropriate Gaussian noise is added to ensure DP. (Song et al.,, 2013; Bassily et al.,, 2014), but with parameters adapted for expert-level privacy on the remainder of the trajectories. For the first filtering part, we use a variant of sparse vector technique for privacy accounting (Dwork and Roth,, 2014). Once this set of trajectory prefixes is obtained, which we call $D^{\Pi}_{stable}$, for the remainder of the trajectories $D^{\Pi}\setminus D^{\Pi}_{stable}$ we run DP-SGD. Our empirical results strongly show that combining steps one and two of the algorithm described above, performs better than either of them individually. The only assumption we make on the underlying offline RL algorithm is that it is gradient-based, so that we can use DP-SGD to privatize its updates.
2 RELATED WORK
Comparison to prior work on DP Offline RL: There have been recent work exploring the intersection of differential privacy (DP) and reinforcement learning (RL) (Wang and Hegde,, 2019; Qiao and Wang,, 2023, 2024; Zhou,, 2022; Liao et al.,, 2023)(Qiao and Wang,, 2023, 2024; Zhou,, 2022) explores the problem either in the tabular setting or in the linear function approximation setting. In either of the cases there is a sufficient statistic (like discrete state-action visit counts or feature covariances to privatize sensitive data) which if made DP, would suffice for the whole algorithm to be DP. In our setting, these algorithms are not applicable because we operate in continuous state space, and with general functional approximation for the RL tasks.
Additionally, Wang and Hegde, (2019) explores the use of differential privacy (DP) in continuous spaces to only protect the value function approximator. Liao et al., (2023) explores RL with linear function approximation under local differential privacy guarantee, a stronger notion of privacy where users contributing data do not trust the central aggregator. For reasons mentioned earlier, these works are also not applicable to our setting.
Furthermore, to the best of our knowledge, expert-level DP has not been dealt with in the context of RL so far. Prior work (cited above) are often limited to exploring trajectory-level privacy, which has finer granularity of privacy protection which makes the setting arguably “easier”. In this work, we adopt an expert-level notion of neighbourhood when a large number of behavioural policies (or experts) are used for data collection.
Background on DP-SGD and Sparse-vector: DP-SGD first introduced in (Song et al.,, 2013; Bassily et al.,, 2014), and adapted for deep learning by Abadi et al., (2016), proposed gradient noise injection to guarantee differentially private model training. Several variants of DP-SGD have been proposed such as McMahan et al., (2017), which allows user-level differentially private training in federated learning scenarios and Kairouz et al., (2021) a private variant of the Follow-The-Regularized-Leader algorithm that compares favourably with amplified DP-SGD (privacy amplification by sampling/shuffling (Abadi et al.,, 2016; Bassily et al.,, 2014)).
Sparse-vector algorithm (Roth and Roughgarden,, 2010; Hardt and Rothblum,, 2010; Dwork and Roth,, 2014) develovep over a sequence of works, allows one to carefully track the prviacy budget across queries, and often provide tighter privacy guarantee as opposed to vanilla composition Dwork and Roth, (2014). At a high-level, the idea is to identify “stable queries”, i.e., the queries that do not change their answers by moving to neighboring data set, and not pay the cost of privacy for answering those queries. In Algorithm 2, we use a modification of this idea to output parts of the expert trajectories that are “stable” without incurring the cost of privacy (in $\varepsilon$).
Relationship with the offline RL literature: As mentioned above, our approach is agnostic to the underlying offline RL algorithm used, and can be combined with off-the-shelf model-based(Kidambi et al.,, 2020; Yu et al.,, 2020, 2021) or model-free (Kumar et al.,, 2019, 2020; Cheng et al.,, 2022) offline RL algorithms. We focus on model-free approaches in our experiments.
3 PRELIMINARIES
We start by introducing the relevant background in offline RL and differential privacy, before formally defining the setting of offline RL with expert-level privacy.
3.1 Offline RL
Offline RL is a data-driven approach to RL that aims to circumvent the prohibitive cost of interactive data collection in many real-world scenarios. Under this paradigm, a static dataset collected by some behaviour policy $\pi_{\beta}$ is used to train another policy $\pi$ without any further interaction with the environment.
Let the environment be represented by a Markov Decision Process (MDP) $\mathcal{M}=\langle S,A,r,P,\rho_{0},\gamma\rangle$, where $S$ is the state space, $A$ is the action space, $r:S\times A\rightarrow\mathbb{R}$ is the (deterministic) reward function, $P:S\times A\to\Delta(S)$ is the transition kernel ($\Delta(\mathcal{X})$ is the set of all probability distributions over $\mathcal{X}$), $\rho_{0}\in\Delta(S)$ is the initial state distribution and $\gamma\in[0,1]$ is the discount factor.
Definition 1 (Offline RL).
[…]
A key concern in offline RL is that the state-action distribution encountered in the offline data to learn a policy $\pi$ might significantly differ from that encountered upon actually executing $\pi$ in the MDP. There is a host of existing offline RL methods that address this challenge in different ways (see e.g. (Kumar et al.,, 2019; Fujimoto et al.,, 2019; Kumar et al.,, 2020; Cheng et al.,, 2022)). We use CQL (Kumar et al.,, 2020) as our offline RL method due to its strong performance across a variety of tasks in prior evaluations.
We consider a scenario where the dataset contains demonstrations collected from a number of different behavioural policies or experts with varying degrees of optimality. We denote the set of behaviour policies by $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. Note that a requirement of expert-level privacy, which we build towards next, can be in tension with the rewards our learned policy can attain. For instance, if most experts are similar, and one which disagrees from them obtains a very high-reward, then there is an inevitable tradeoff between expert-level privacy and utility.
3.2 Differential Privacy
Differential privacy (DP) was first introduced in Dwork et al., (2006) to quantify the loss of privacy associated with data released drawn from any statistical database. Informally, a procedure satisfies DP if the distribution of its outputs on two datasets which differ only in one record is very similar. This notion is formally defined below.
Definition 2 ($(\varepsilon,\delta)$-DP).
For positive real numbers $\varepsilon>0$ and $0\leq\delta\leq 1$, an algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP iff, for any two neighbouring datasets $D,D^{\prime}\in\mathcal{D}^{*}$ ($\mathcal{D}$ is the space of all data records), and $S\subset Range(\mathcal{A})$:
$Pr(\mathcal{A}(D)\in S)\leq e^{\varepsilon}Pr(\mathcal{A}(D^{\prime})\in S)+\delta$ (1)
where $Range(\mathcal{A})$ is the set of all possible outputs of $\mathcal{A}$.
Definition 2 guarantees that the probability of seeing a specific output on any two neighbouring datasets can differ at most by a factor of $e^{\varepsilon}$, with an additional relaxation of $\delta$. In most ML applications reasonable privacy guarantees are provided by $\varepsilon\leq 10$ (Ponomareva et al.,, 2023) and $\delta=n^{-1.1}$, where $n$ is the number of individuals in the dataset.
One can instantiate Definition 2 with different of neighbourhood relations. Next, we describe the neighbourhood relations underlying our setting of expert-level privacy and the more commonly studied trajectory-level privacy.
Definition 3 (Expert- and Trajectory-level privacy).
Let $\mathcal{D}$ be the domain of all valid trajectories possible in the underlying MDP. Let $D^{\Pi}\in\mathcal{D}^{*}$ be a data set of trajectories generated by a set of experts $\Pi$. We say $D^{\Pi},D^{\Pi^{\prime}}\in\mathcal{D}^{*}$ to be expert-level neighbours if the set of experts $|\Pi\Delta\Pi^{\prime}|=1$, where $\Delta$ is the set difference. Alternatively, if $D,D^{\prime}\in\mathcal{D}^{*}$ are sets of trajectories and $|D\Delta D^{\prime}|=1$, then we say $D$ and $D^{\prime}$ are trajectory-level neighbours.
We note that the key difference between expert and trajectory level privacy notions, as defined above is that removing an expert $\pi$ can remove up to $|D^{\pi}|$ trajectories from the overall dataset, where $D^{\pi}$ is the set of trajectories contributed by $\pi$. This is a much larger change than allowed in trajectory-level privacy, and a naïve treatment would result in paying an additional cost due to group privacy (Dwork and Roth,, 2014), scaling linearly in the number of trajectories each expert contributes. At the same time, as we argue below, via a set of settings, expert-level privacy is often a critically required notion of privacy protection. We also note that we do not include the MDP dynamics and rewards, $P$ and $r$ in the private information we protect, since that would require incorporating changes to the MDP in our neighbourhood relation too.
Example (Personalized healthcare interventions): Many works (Saria,, 2018; Gottesman et al.,, 2019; Tang et al.,, 2022) study the promise of offline RL for learning individual treatment recommendations, using existing clinical data, since real-world data interventions in the form of clinical trials are extremely expensive and limited in their power for learning personalized treatments. Here, the actions comprise of the units of medication administered to an intensive-care patient, now a well-studied setting in the case of sepsis care (Oberst and Sontag,, 2019). States are outcomes of tests as well as relevant information from the patient’s health record including demographic information, recent measaurements of relevant variables such as heart rate, blood pressure etc. The experts correspond to doctors or clinics who choose the observed treatment regimes in the offline samples, and we would like to learn high-performing policies without violating their privacy. Rewards correspond to long-term health outcomes such as recovery time.
As evident in the example above, having expert-level privacy can reduce the barrier for the decision makers to make the decision trajectories available for learning, without worrying about leakage of proprietary information or regulatory/litigation concerns. We note that expert-level privacy differs from the privacy of a patient in the healthcare domains, or broader privacy notions also including the transitions and rewards of the MDP, since we focus on the feasibility of learning from the experts’ choices via offline RL.
3.3 Problem Setup
Dataset.
Let $\Pi$ denote a set of $m$ private policies, corresponding to $m$ experts, $\Pi=\{\pi_{1},\dots,\pi_{m}\}$. $D^{\Pi}=\{D^{\pi_{1}},\dots,D^{\pi_{m}}\}$ is the aggregated non-private dataset generated by the interaction of the $m$ private policies with the environment. Each $D^{\pi_{i}}$ is a set of $N$ trajectories $\tau=(s_{1},a_{2},s_{2},a_{2},\dots,s_{L},a_{L},s_{L+1})$ of length $L$ obtained by logging the interactions of policy $\pi_{i}$ with the MDP $\mathcal{M}=\langle S,A,r,P,\rho_{0},\gamma\rangle$, where $s_{1}\sim\rho_{0}$, $a_{h}\sim\pi_{i}(\cdot|s_{h})$ and $s_{h+1}\sim P(\cdot|s_{h},a_{h})$. For notational simplicity we omit rewards throughout the paper as they do not play a role in our privatization approach and we can just assume they can be computed as needed for a given state-action pair.
We denote by $|\tau|$ the length of the trajectory as the number of full state-action pairs it consists of. Further, we use $\tau_{h:h^{\prime}}=(s_{h},a_{h},\dots s_{h},a_{h},s_{h^{\prime}+1})$ to denote the sub-trajectory of $\tau$ starting from the $h$-th timestep until the $h^{\prime}$-th, including the trailing next state $s_{h^{\prime}+1}$.
Goal.
The goal is to learn a policy using an offline RL algorithm with expert-level DP guarantees on $D^{\Pi}$, thus ensuring that the policy learnt in this manner is not much different from a policy learnt using $D^{\Pi^{\prime}}$. We make the following assumptions.
Assumption 1.
The action space $A$ is discrete.
This is a fairly mild assumption satisfied in many natural settings. The assumption arises in the first-phase of identifying stable trajectory prefixes for our algorithm, and extending this to continuous action spaces is an interesting direction for future work.
Assumption 2.
We can query the private expert policies, i.e., evaluate $\pi_{i}(a|s)$ for a given $i,s$ and $a$.
Again, this query access only comes up in computing the stability of the trajectory prefix, and can likely be relaxed to using behavior-cloned versions of the experts with a slightly worse utility bound.
Assumption 3.
Given $p_{\min}\in(0,1/|A|)$, define
$\Pi_{p_{\min}}=\{\pi\colon\inf_{s\in S}\min_{a\in A}\pi(a|s)\geq p_{\min}\},$
as the set of all stochastic policies with minimum action probability at least $p_{\min}$. Then all expert policies are in this set, that is, $\Pi\subseteq\Pi_{p_{\min}}$ and our privacy guarantee only considers policies in $\Pi_{p_{\min}}$.
This last assumption is purely technical and an artifact of our analysis. Without a guaranteed $p_{\min}$ the ratio of action probabilities between two neighboring databases can be infinite, causing a problem in our proofs. We look forward to removing this assumption in future work.
4 ALGORITHM
[…]
Algorithm 1 PrefixQuery ($\mathcal{A}_{PQ}$): Tests if count of experts expected to execute a trajectory is large enough
1:Expert policies $\Pi$, trajectory $\tau=(s_{1},a_{1},s_{2},a_{2},\dots)$, stability threshold $\hat{\theta}$, $\varepsilon$
2:Compute query count as
3:$count_{\tau}(\Pi)\leftarrow\sum_{\pi\in\Pi}\prod_{j=1}^{|\tau|}\pi(a_{j}|s_{j})$
4:if $count_{\tau}(\Pi)+Lap(\frac{4}{\varepsilon})>\hat{\theta}$ then
5:     Return $\top$
6:else
7:     Return $\perp$
Algorithm 2 DataRelease ($\mathcal{A}_{DR})$: Releasing public dataset after privatisation
1:Expert dataset $D^{\Pi}$, expert policies $\Pi$, $\varepsilon_{1}$, $\delta_{1}$, min action-selection probability $p_{min}$, unstable query cutoff $T\leq N\cdot m$, trajectory length $L$
2:$\varepsilon^{\prime}\leftarrow\frac{\varepsilon_{1}}{\sqrt{32T\log(2/\delta_{1})}},\quad\delta^{\prime}\leftarrow\frac{\delta_{1}}{(2TL)}$
3:$c_{min}\leftarrow\frac{e^{\varepsilon^{\prime}}}{(e^{\varepsilon^{\prime}}-1)},\quad\theta\leftarrow\frac{c_{min}}{p_{min}},\quad D^{\Pi}_{stable}\leftarrow\{\}$
4:Reshuffle the order of $D^{\Pi}$
5:for $c=1,\dots,T$ do
6:     $\tau\leftarrow$ next trajectory from $D^{\Pi}$
7:     $\hat{\theta}\leftarrow\theta+\frac{4}{\varepsilon^{\prime}}\log(1/\delta^{\prime})+Lap(\frac{2}{\varepsilon^{\prime}})$
8:     for $i=1$ to $L$ do
9:         $r\leftarrow\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$triangleright Algorithm 1
10:         if $r=\top$ then
11:              if $i=L$ then
12:                  $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i}\}$
13:         else
14:              if $i\neq 1$ then
15:                  $D^{\Pi}_{stable}\leftarrow D^{\Pi}_{stable}\cup\{\tau_{1:i-1}\}$
16:              break
17:return $D^{\Pi}_{stable}$
Algorithm 2 uses Algorithm 1 as a subroutine, traversing up to $T$ trajectories as long as the prefix remains stable. The dataset $D^{\Pi}_{stable}$ returned by Algorithm 2 is composed of these stable trajectory prefixes. Any trajectories after the first $T$ for a suitably chosen parameter $T$ are not included in $D^{\Pi}_{stable}$ to respect the privacy budget. The discarded tails of each trajectory (including the trajectories that Algorithm 2 did not process) are collected in another dataset:
$D^{\Pi}_{unst}=\{\tau_{k+1:L}\mbox{ if }\tau_{1:k}\in D^{\Pi}_{stable}\mbox{ for $k\geq 1$ or }\tau,~{}~{}\tau:D^{\Pi}\}$
Consider any parameterized gradient-based offline RL algorithm $\mathcal{H}_{\phi}$. Training this algorithm consists of using $(s,a,r,s^{\prime})$ tuples collected in the offline dataset to update the parameters (by splitting the trajectories into into individual transitions). Having split the dataset into stable and unstable trajectories, we can now leverage the stability of $D^{\Pi}_{stable}$ to accumulate the gradient updates of $\mathcal{H}_{\phi}$ on transitions drawn from this dataset without any noise addition. The full algorithm, detailing the training using a mix of noisy gradient updates for transitions from $D^{\Pi}_{unst}$ and noiseless gradient updates for transitions from $D^{\Pi}_{stable}$ is given in Algorithm 3.
Algorithm 3 Selective DP-SGD for Offline RL
1:Stable dataset $D^{\Pi}_{stable}$, Unstable dataset $D^{\Pi}_{unst}$, Sampling probability $p$, Gradient-based offline RL algorithm $\mathcal{H}_{\phi}$, $\varepsilon_{2},\delta_{2}$, $N$ training steps, $B$ batch size
2:Initialize $\phi_{0}$
3:for $n=1$ to $N$ do
4:     $b\sim Bernoulli(p)$
5:     if $b=1$ then $\triangleright$ w.p. $p$
6:         Sample batch of transitions $\mathcal{B}\sim D^{\Pi}_{unst}$
7:         $\phi_{n+1}\leftarrow$ DP-SGD($\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N$)
8:     $\triangleright$ w.p. $1-p$
9:         Sample batch of transitions $\mathcal{B}\sim D^{\Pi}_{stable}$
10:         $\phi_{n+1}\leftarrow$ SGD($\phi_{n},\mathcal{B}$)
11:return $\mathcal{H}_{\phi_{N}}$
The sampling probability $p$ controls the expected fraction of timesteps sampled from the unstable dataset $D^{\Pi}_{unst}$. The DP-SGD updates of the form DP-SGD($\phi_{n},\mathcal{B},\varepsilon_{2},\delta_{2},N$) in Algorithm 3 refer to noisy gradient updates with additive noise scaled as to ensure $(\varepsilon_{2},\delta_{2})$ expert-level DP. To do so we add the modifications to the standard DP-SGD algorithm (Abadi et al.,, 2016):
  • During gradient updates, we ensure that an expert contributes at most one transition to a batch.
  • To account for privacy amplification while sampling batches from the dataset, we use the use the fraction of experts from which we sample. Hence, subsampling coefficient $q$ for DP-SGD analysis is $q=p\cdot\frac{|B|}{m}$, where $m$ is the number of experts.
A detailed description of the modified expert-level DP-SGD, which we also employ as a baseline in our empirical work in Section 6 is provided in Appendix B. The privacy analysis for the overall pipeline combining Algorithms 2 and 3 is discussed in Section 5.
Sub-optimality guarantees for the learned policy: While we show in the following section that the Algorithm 3 provides expert-level privacy, we do not provide explicit bounds on the suboptimality of the returned policy. Typical offline RL literature shows that using pessimistic approaches, the learned policy is competitive with any other policy whose state-action distribution is well-covered by the offline data distribution (Xie et al.,, 2021; Cheng et al.,, 2022). However, our training distribution undergoes two changes from the apriori offline data distribution. First is that the stable prefixes might be distributionally rather different from the entire state-action distribution, and potentially emphasize the states and actions near the start of a trajectory, as shorter prefixes are more likely to be stable. Second, the gradient clipping in DP-SGD adds an additional bias. Understanding the effect of latter with arbitrary function approximation is beyond the scope of this paper. We do, however, study the effects of distributional biases from adding varying amounts of the stable prefixes in our experiments, and find that they are typically very beneficial, particularly in tasks where the optimal behaviour does not change significantly over time.
5 PRIVACY ANALYSIS
For the privacy analysis of our proposed algorithm, we adopt notations for the MDP described in Section 3.1 and problem setup described in Section 3.3. We assume for the benefit of the analysis, that all such trajectories are of the same length $L$ (number of actions taken before the episode ends). Practically this can be replaced by the length of the longest trajectory for a worst-case privacy analysis.
Algorithm 2 can be thought of as $T$ iterations of a combination of two algorithms $\mathcal{A}_{1},\mathcal{A}_{2}$:
  • $\mathcal{A}_{1}$ samples a random expert $\pi\sim\mathrm{Unif}(\Pi)$ and a trajectory $\tau\sim\pi$ and converts it into $L$ queries, $Q^{\tau}=\{count_{\tau_{1}}(.),...,count_{\tau_{L}}(.)\}$.
  • $\mathcal{A}_{2}$ is invoked on $\Pi$ and the query set $Q^{\tau}$, running an iteration of the Sparse Vector Technique (Dwork and Roth, (2014)). $\mathcal{A}_{2}$ treats $\Pi$ as the private database, $Q^{\tau}$ as the query sequence, and $\hat{\theta}$ as the threshold for the Sparse Vector Algorithm. $\mathcal{A}_{2}(\Pi,Q^{\tau})$ outputs $r\in\{\top,\bot\}^{k+1}$ such that $r[i]=\top\;\forall\;i\leq k$ and $r[k+1]=\bot$, where $k\leq L$ is the length of the stable prefix. Note that $r[i]=\mathcal{A}_{PQ}(\Pi,\tau_{1:i},\hat{\theta},\varepsilon^{\prime})$.
  • $\mathcal{A}_{1}$, upon receiving $r$, releases the prefix $\tau_{1:k}=(s_{1},a_{1},\dots,s_{k},a_{k},s_{k+1})$.
For a fixed stream of queries $Q^{\tau}$, Algorithm $\mathcal{A}_{2}$ is $(\varepsilon^{\prime},0)$-DP since it just invokes one round of the Sparse Vector Algorithm (Dwork and Roth, (2014)). For a given trajectory $\tau$, $\mathcal{A}_{2}$ identifies the longest prefix which is stable, or in other words, has enough experts which are likely to sample this prefix. Note that, the privacy cost $\varepsilon^{\prime}$ for identifying this prefix is independent of the length of the trajectory.
However, $Q^{\tau}$ itself is a function of the set of experts $\Pi$. Despite this, in the following analysis, we show that $\mathcal{A}=\mathcal{A}_{1}\circ\mathcal{A}_{2}$ is also differentially private (due to the carefully chosen threshold in Algorithm 2 Line 6). We then compose the privacy guarantees over $T$ trajectories, since $\mathcal{A}_{DR}$ is simply $\mathcal{A}$ applied $T$ times.
We now state two key lemmas for proving that $\mathcal{A}$ is differentially private. For a given trajectory prefix $\tau_{k}$ of length $k$, let $E_{\tau}$ denote the event that $\mathcal{A}_{1}$ samples any trajectory $\tau^{\prime}$ containing $\tau_{k}$ as a prefix; that is, $\tau^{\prime}_{1:k}=\tau_{k}$.
Lemma 5.1.
For a trajectory prefix $\tau_{k}$, and neighbouring expert sets $\Pi,\Pi^{\prime}$, if $count_{\tau_{k}}(\Pi)\geq c_{min}$ then,
$\Pr(E_{\tau_{k}}|\Pi)\leq e^{\varepsilon^{\prime}}\Pr(E_{\tau_{k}}|\Pi^{\prime})$ (2)
where $c_{min}$ and $\varepsilon^{\prime}$ are as defined in Algorithm 2.
This lemma states that if the probability of seeing a trajectory prefix $\tau_{k}$ across all experts is large enough, then the probability that $\mathcal{A}_{1}$ samples any trajectory with prefix $\tau_{k}$ is stable under changing one of the experts. The full proof is available in Appendix A.
Lemma 5.2.
For any trajectory prefix $\tau_{k}$, and expert set $\Pi$ such that $count_{\tau_{k}}(\Pi)<\theta$ the following holds:
$\Pr(\mathcal{A}_{PQ}(\Pi,\tau_{k},\hat{\theta},\varepsilon^{\prime})=\top)<\delta^{\prime}$ (3)
where $\varepsilon^{\prime},\delta^{\prime},\theta,\hat{\theta}$ are as defined in Algorithm 2.
Informally, this lemma states that if the total probability of seeing a trajectory prefix $\tau_{k}$ across all experts is not large enough, then the probability that it is labelled as stable by Algorithm 1 is less than $\delta^{\prime}$. The full proof is described in Appendix A.
Theorem 5.3.
Under Assumptions 1-3, $\mathcal{A}$ is $(2\varepsilon^{\prime},\frac{\delta_{1}}{2T})$-DP, where $\varepsilon^{\prime},\delta_{1}$ are defined in Algorithm 2.
Proof Sketch: The main result used to prove this theorem is the following:
$\Pr(\mathcal{A}(\Pi)=\tau)\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau)+\Pr(E_{\tau}|\Pi)\cdot\delta^{\prime}$ (4)
which is obtained from Lemma 5.4 that we give below.
Let $\mathcal{T}$ denote the set of all possible trajectory prefixes using $S,A$. To prove that $\mathcal{A}$ is $(2\varepsilon^{\prime},\delta_{1}/(2T))$-differentially private, we need to prove the following for any pair of neighbouring $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$:
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}$
$\leq\frac{\delta_{1}}{2T}.$
The above can be proved using (4), and we defer the full proof to Appendix A.
The following lemma gives the key result used to prove Theorem 5.3.
Lemma 5.4.
Fix a trajectory prefix $\tau_{k}$ and a pair of neighboring expert sets $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$. If $count_{\tau_{k}}(\Pi)<c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)=\tau_{k})<\Pr(E_{\tau_{k}}|\Pi)\cdot\delta^{\prime}.$ (5)
If $count_{\tau_{k}}(\Pi)\geq c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)={\tau_{k}})\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})={\tau_{k}})+\Pr(E_{\tau_{k}}|\Pi)\cdot\delta^{\prime}$ (6)
Proof Sketch: The proof of this lemma follows from Lemmas 5.1 and 5.2. Let $r=\top\cdots\top\bot\in\{\top,\perp\}^{k+1}$ and $\gamma^{(L)}=\top^{L}$. When $k<L$ we have
$\Pr(\mathcal{A}(\Pi)=\tau_{k})$
$=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$
where $\tau_{k}\cdot a$ denotes the trajectory prefix $\tau_{k}$ followed by action $a$. And when $k=L$,
$\Pr(\mathcal{A}(\Pi)=\tau_{k})=\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau}_{k})=\gamma^{(L)})\Pr(E_{\tau_{k}}|\Pi)$
(a) LunarLander
(b) Acrobot
(c) CartPole
(d) HIV Treatment
Figure 2: Expert return histograms for experts trained on heterogeneous LunarLander, Acrobot, CartPole and HIV Treatment environments. We use a mix of experts that shows a wide range of performance on the test environment. For HIV Treatment, the return values are normalised by $61e$ 6.
Let’s consider the case $k=L$. The main idea behind this proof is that if $count_{\tau_{k}}(\Pi)$ is small enough, then $\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau}_{k})=\gamma^{(L)})$ is less than $\delta^{\prime}$ (Lemma 5.2) and we get (5). On the other hand, if $count_{\tau_{k}}(\Pi)$ is greater than $c_{min}$, then $\Pr(E_{\tau_{k}}|\Pi)$ is bounded by Lemma 5.1. Using that $\mathcal{A}_{2}$ is always $\varepsilon^{\prime}$-DP when the queries are fixed we get (6). The case of $k<L$ is less straightforward, but can also be analyzed by considering two cases for each $a\in A$, $count_{\tau_{k}\cdot a}(\Pi)\geq c_{min}$ or $count_{\tau_{k}\cdot a}(\Pi)<c_{min}$. The full proof is present in Appendix A. ∎
Theorem 5.5.
Under Assumptions 1-3, Algorithm 2 is $(\varepsilon_{1},\delta_{1})$-DP.
Proof.
Algorithm 2 simply executes Algorithm $\mathcal{A}$, $T$ times. Since $\mathcal{A}$ is $(\varepsilon_{1}/\sqrt{8T\log(2/\delta_{1})},\delta_{1}/(2T))$-DP as shown in Theorem 5.3, using advanced composition (Corollary 3.21 in Dwork and Roth, (2014)), Algorithm 2 can be shown to be $(\varepsilon_{1},\delta_{1})$-DP. ∎
Using all the above, we derive the following expert-level privacy guarantee for the overall training pipeline.
Theorem 5.6.
Algorithm 3 is $(\varepsilon,\delta)$-DP, where:
$\varepsilon=\varepsilon_{1}+\varepsilon_{2}\quad\text{and}\quad\delta=\delta_{1}+\delta_{2}$ (7)
Proof.
Shown trivially by sequential composition of privacy guarantees for the release of $D^{\Pi}_{stable}$ and DP-SGD training on samples from $D^{\Pi}_{unst}$. ∎
6 EXPERIMENTS
Ensuring reasonable levels of privacy requires significant amounts of hyperparameter tuning and a large number of data points, and typically comes with a non-trivial performance impact [^3] In supervised learning, for example, the accuracy on ImageNet training can go from $>75\%$ to $\sim 47.9\%$ in the private case (Kurakin et al.,, 2022). Intuitively, with the demand for expert-level privacy in our setting, we risk severe degradation of utility since each “record” (each expert’s data) forms a much larger fraction of the overall dataset. It is thus imperative that we create a dataset which consists of diverse data collected from a large number of different experts. We describe our data generation procedure to ensure this below.
6.1 Dataset Generation
We train $3000$ experts each on the Cartpole, Acrobot, LunarLander and HIV Treatment environments. Cartpole, Acrobot and LunarLander are from the Gymnasium package (Towers et al.,, 2023). The HIV Treatment simulator is based on the implementation by Geramifard et al., (2015) of the model described in Ernst et al., (2006). Experts are trained on variations of the default environment, created by modifying the values of key parameters (e.g. gravity in LunarLander environment). The trained experts are then used to generate demonstrations on a default setting of the environment, collectively forming the aggregated dataset for offline RL. This training methodology ensures that the experts show varied (often sub-optimal) behaviours on the default environment. As shown in Figure 2, the experts’ return-histograms indicate that the mix of experts we create shows significant diversity in the policies learnt. Further details regarding the environment settings used to train the expert policies are avilable in Appendix B. We acknowledge that due to Assumption 1, we do not consider continuous control domains here and focus on discrete action benchmarks.
6.2 Training Setup
While our algorithm allows us to use any off-the-shelf gradient-based offline RL algorithm, we use Conservative Q-Learning (CQL) (Kumar et al.,, 2020) algorithm for our experimental work. More specifically, we use the DQN version of CQL, since we are dealing with discrete action spaces. To represent the Q-value function, we use a neural network with 2 hidden layers of 256 units each. During training, we perform a grid search to find the optimal set of hyper-parameters (learning rate, sampling probability $p$, DP-SGD noise) for each environment. The batch size is fixed to 256 for Acrobot, LunarLander and Cartpole, and to 64 for HIV Treatment.
The maximum trajectory length is fixed to 200 for Acrobot and CartPole in the collected demonstrations. While evaluating the final learned policy, the maximum episode length is again set as 200 for Acrobot, we increase this to 1000 for Cartpole for increased difficulty in an otherwise “simple” environment. We allow all DP-SGD training runs to progress until the privacy budget is used up completely and report statistics of episode returns averaged over $10$ evaluation runs at the end of training.
We keep the privacy budget constant to $\varepsilon=10$, and $\delta=1/m$, where $m$ is the number of experts.
6.3 Baselines
We compare our method against DP-SGD adapted for expert-level privacy. A brief discussion of the changes we made are presented in Section 4. A full discussion, along with analysis, is presented in the appendix for reference. We keep the underlying RL algorithm as CQL for fair comparison. Since there are no prior methods for the expert-level privacy setting, we do not have other readily available baselines.
We also compare against a non-private baseline, which ignores privacy concerns and trains over the whole dataset using vanilla CQL. This may be taken as an upper-bound on the performance that we hope to achieve in the presence of privacy concerns.
6.4 Results
(a) LunarLander
(b) Acrobot
(c) CartPole
(d) HIV Treatment
Figure 3: Our method performs favourably compared to expert-level DP-SGD on multiple test environments. The average episodic return over $10$ evaluation runs at the end of training are reported along with respective confidence intervals. For HIV Treatment, the return values are normalised by $61e$ 6.
(a) LunarLander
(b) Acrobot
(c) CartPole
Figure 4: Performance of our method and DP-SGD for different values of $\varepsilon$. We see consistent improvements of our method over DP-SGD in the higher $\varepsilon$ regions.
Our experiments show that our algorithm which leverages expert consensus to select stable trajectory prefixes prior to training CQL, shows superior performance to training CQL by merely using an adapted expert-level DP-SGD algorithm. Our method is able to recover much of the performance loss due to DP-SGD which is evident by our results being very close to the non-private training results (Figure 3).
We also provide the values of sampling probability that led to the best performance for each environment in Table 1. Interestingly, for Acrobot and HIV Treatment environments, just $D_{stable}^{\Pi}$ was enough to get the optimal performance and we did not need to pay any privacy cost of DP-SGD during model training.
Table 1: Best choices for probability of sampling from $\mathcal{D}^{\Pi}_{stable}$ ie. $1-p$ for Algorithm 3 on test environments
Environment Sampling probability
LunarLander 0.2
Acrobot 1.0
CartPole 0.1
HIV Treatment 1.0
Figure 4 compares the performance of our method with DP-SGD across different values of $\varepsilon$. All the results are averaged over 3 runs and $95\%$ confidence intervals are also reported. Note that our method degenerates to DP-SGD if we take $\varepsilon_{1}=0,\varepsilon_{2}=\varepsilon$ for any given $\varepsilon$, which is what we do in the lower $\varepsilon$-region ($\varepsilon\leq 5$ for Lunar Lander and CartPole, and $\varepsilon\leq 2.5$ for Acrobot) where the privacy budget is not enough to get sufficient stable data for training.
For our method, for higher values of $\varepsilon$, we take $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4}$ for LunarLander and CartPole. For Acrobot, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0$, that is, we only train on $D_{stable}^{\Pi}$ for Acrobot. Similarly, we split $\delta$ as $\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$.
7 CONCLUSION AND FUTURE WORK
In this work, we initiate the study of offline RL with expert-level privacy. We describe the novel challenges of the setting, and provide theoretically sound algorithms which are evaluated in proof-of-concept evaluation. In future work, we would like to relax Assumptions 1-3, which limit the applicability of our results. We expect replacing the queries to the experts with behaviour cloned versions to be less challenging than eliminating the other two assumptions. Understanding the utility implications of Algorithm 3, despite the influence of our privatization scheme on the training distribution is another important avenue for future work. Finally, while our empirical evaluation convincingly establishes the promise of this approach, we leave a larger evaluation across diverse and larger-scale RL benchmarks for future studies.
Appendix A Theorem Proofs
A.1 Proof of Lemma 5.1
See 5.1
Proof.
Recall $E_{\tau_{k}}$ is the event a trajectory with prefix $\tau_{k}$ is generated. Assuming $\tau_{k}=s_{1},a_{1},s_{2},\ldots,a_{k},s_{k+1}$, the probability that a fixed expert $\pi$ generates such a prefix when interacting with the MDP is $\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}\pi(a_{j}|s_{j})P(s_{j+1}|s_{j},a_{j})\right)$. Since the expert is chosen uniformly at random, we have:
$\Pr(E_{\tau_{k}}|\Pi)$ $=\frac{1}{m}\sum_{i=1}^{m}\left[\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}\pi_{i}(a_{j}|s_{j})P(s_{j+1}|s_{j},a_{j})\right)\right]$ (8)
$=\frac{1}{m}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)\sum_{i=1}^{m}\left(\Pi_{j=1}^{k}\pi_{i}(a_{j}|s_{j})\right)$ (9)
$=\frac{1}{m}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi)\enspace.$ (10)
Similarly,
$\Pr(E_{\tau_{k}}|\Pi^{\prime})=\frac{1}{m^{\prime}}\rho_{0}(s_{1})\left(\Pi_{j=1}^{k}P(s_{j+1}|s_{j},a_{j})\right)count_{\tau_{k}}(\Pi^{\prime})\enspace.$ (11)
Where $m^{\prime}=|\Pi^{\prime}|$. Since $|\Pi\Delta\Pi^{\prime}|=1$,
$|count_{\tau_{k}}(\Pi)-count_{\tau_{k}}(\Pi^{\prime})|\leq 1\quad\text{and}\quad|m-m^{\prime}|=1\enspace.$ (12)
This gives us the following,
$\frac{\Pr(E_{\tau_{k}}|\Pi)}{\Pr(E_{\tau_{k}}|\Pi^{\prime})}$ $=\frac{count_{\tau_{k}}(\Pi)\times m^{\prime}}{count_{\tau_{k}}(\Pi^{\prime})\times m}$ (13)
$\leq\max\left(\frac{count_{\tau_{k}}(\Pi)(m+1)}{(count_{\tau_{k}}(\Pi)+1)m},\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m},\frac{m-1}{m}\right)$ (14)
$=\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)\enspace.$ (15)
Note that, since $m+1>count_{\tau_{k}}(\Pi)$,
$\frac{m+1}{m}<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}$ (16)
$\max\left(\frac{count_{\tau_{k}}(\Pi)(m-1)}{(count_{\tau_{k}}(\Pi)-1)m},\frac{m+1}{m}\right)<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\enspace.$ (17)
This allows us to bound the ratio of probabilities of $E_{\tau_{k}}$ for $\Pi$ and $\Pi^{\prime}$ as
$\frac{\Pr(E_{\tau_{k}}|\Pi)}{\Pr(E_{\tau_{k}}|\Pi^{\prime})}<\frac{count_{\tau_{k}}(\Pi)}{count_{\tau_{k}}(\Pi)-1}\leq e^{\varepsilon^{\prime}}\enspace,$ (18)
where the last inequality uses $count_{\tau}(\Pi)\geq c_{min}=e^{\varepsilon^{\prime}}/(e^{\varepsilon^{\prime}}-1)$. ∎
A.2 Proof of Lemma 5.2
See 5.2
Proof.
$\Pr(\mathcal{A}_{PQ}(\Pi,\tau,\hat{\theta},\varepsilon^{\prime})=\top)$ $=\Pr\left[count_{\tau_{k}}(\Pi)+Lap(4/\varepsilon^{\prime})>\theta+(4/\varepsilon^{\prime})\log(1/\delta^{\prime})+Lap(2/\varepsilon^{\prime})\right]$ (19)
$<\Pr\left[\theta+Lap(4/\varepsilon^{\prime})>\theta+(4/\varepsilon^{\prime})\log(1/\delta^{\prime})+Lap(2/\varepsilon^{\prime})\right]$ (20)
$=\Pr\left[Lap(4/\varepsilon^{\prime})>(4/\varepsilon^{\prime})\log(1/\delta^{\prime})+Lap(2/\varepsilon^{\prime})\right]$ (21)
Consider $u\sim Lap(4/\varepsilon^{\prime})$ and $v\sim Lap(2/\varepsilon^{\prime})$. For a fixed $v$, we have,
$\Pr\left[u\geq(4/\varepsilon^{\prime})\log(1/\delta^{\prime})+v\right]\leq\frac{\delta^{\prime}}{2}e^{-v(\varepsilon^{\prime}/4)}$ (22)
[Using Laplace Tail Bound]
Accounting for the randomness of $v$, from (21), (22) and the pdf of the Laplace distribution, we have,
[…]
$<\delta^{\prime}$ (24)
This gives the statement of the lemma.
A.3 Proof of Lemma 5.4
See 5.4
Proof.
We split this proof across 3 lemmas. Lemma A.1 proves this bound for the case $count_{\tau_{k}}(\Pi)<c_{min}$ and $k<L$. Lemma A.2 handles the case $count_{\tau_{k}}(\Pi)\geq c_{min}$ and $k<L$. Finally, Lemma A.3 proves this for $k=L$. Combining these 3 gives the statement of Lemma 5.4. ∎
Lemma A.1.
For any trajectory prefix $\tau_{k},k<L$, and neighboring expert sets $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$, if $count_{\tau_{k}}(\Pi)<c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)=\tau_{k})<\Pr(E_{\tau_{k}}|\Pi)\delta^{\prime}$
where $\delta^{\prime}$ is as defined in Algorithm 2.
Proof.
Let $r=\top\cdots\top\bot\in\{\top,\perp\}^{k+1}$
$\Pr(\mathcal{A}(\Pi)=\tau_{k})=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$ (25)
where $\tau_{k}\cdot a$ denotes the trajectory prefix $\tau_{k}$ followed by action $a$.
[…]
$\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\leq\Pr(\mathcal{A}_{PQ}(\Pi,\tau_{k},\hat{\theta},\varepsilon^{\prime})=\top)$ (26)
afor-$\forall a$, Using $count_{\tau}(\Pi)<c_{min}<\theta$, (since $\theta=c_{min}/p_{min}$), Lemma 5.2, and (26),
$\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\leq\delta^{\prime}$ (27)
Combining (25) and (27), we get,
$\Pr(\mathcal{A}(\Pi)=\tau_{k})<\delta^{\prime}\sum_{a\in A}\Pr(E_{\tau_{k}\cdot a}|\Pi)=\delta^{\prime}\Pr(E_{\tau_{k}}|\Pi)$ (28)
Lemma A.2.
For any trajectory prefix $\tau_{k}=(s_{1},a_{1},...,s_{k},a_{k},s_{k+1}),k<L$, and neighboring expert sets $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$, if $count_{\tau_{k}}(\Pi)\geq c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)=\tau_{k})\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau_{k})+\Pr(E_{\tau_{k}}|\Pi)\delta^{\prime}$
where $\varepsilon^{\prime},\delta^{\prime}$ are as defined in Algorithm 2.
Proof.
Let $r=\top\cdots\top\bot\in\{\top,\perp\}^{k+1}$.
$\Pr(\mathcal{A}(\Pi)=\tau)=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau\cdot a})=r^{k+1})\Pr(E_{\tau\cdot a}|\Pi)\right]$ (29)
Let $A^{\prime}=\{a:a\in A,count_{\tau_{k}\cdot a}(\Pi)\geq c_{min}\},A^{\prime\prime}=A\setminus A^{\prime}$.
$\forall a\in A^{\prime\prime}$,
$count_{\tau_{k}\cdot a}(\Pi)=\sum_{i=1}^{m}[(\Pi_{j=1}^{k}\pi_{i}(a_{j}|s_{j}))\times\pi_{i}(a|s_{k+1})]\geq p_{min}\times\sum_{i=1}^{m}[\Pi_{j=1}^{k}\pi_{i}(a_{j}|s_{j})]=p_{min}\times count_{\tau_{k}}(\Pi)$ (30)
$p_{min}\times$ $count_{\tau_{k}}(\Pi)\leq count_{\tau_{k}\cdot a}(\Pi)<c_{min}$ (31)
$count_{\tau_{k}}(\Pi)<\frac{c_{min}}{p_{min}}=\theta$ (32)
The output of $\mathcal{A}_{2}$ is the output of a sequence of calls to $\mathcal{A}_{PQ}$ (Algorithm 1). The probability that $\mathcal{A}_{2}$ outputs the sequence $r$ can be upper-bounded by the probability of getting $\top$ on the $k^{th}$ call to $\mathcal{A}_{PQ}$. Using this fact along with (32) and Lemma 5.2, we can say,
$\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})$ $=r)\leq\Pr(\mathcal{A}_{PQ}(\Pi,\tau_{k},\hat{\theta},\varepsilon^{\prime})=\top)<\delta^{\prime}$ (33)
$\Pr(\mathcal{A}(\Pi)=\tau_{k})$ $=\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$ (34)
$=\sum_{a\in A^{\prime}}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$
$\qquad+\sum_{a\in A^{\prime\prime}}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$ (35)
$=\sum_{a\in A^{\prime}}\left[\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi)\right]$
[…]
$\leq e^{2\varepsilon^{\prime}}\sum_{a\in A^{\prime}}\left[\Pr(\mathcal{A}_{2}(\Pi^{\prime},Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi^{\prime})\right]$
[…]
$\leq e^{2\varepsilon^{\prime}}\sum_{a\in A}\left[\Pr(\mathcal{A}_{2}(\Pi^{\prime},Q^{\tau_{k}\cdot a})=r)\Pr(E_{\tau_{k}\cdot a}|\Pi^{\prime})\right]+\delta^{\prime}\Pr(E_{\tau_{k}}|\Pi)$ (38)
$\Pr(\mathcal{A}(\Pi)=\tau_{k})$ $\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau_{k})+\Pr(E_{\tau_{k}}|\Pi)\delta^{\prime}$ (39)
Lemma A.3.
For any trajectory prefix $\tau_{L}$, where $L$ is the maximum length of a trajectory, and a pair of neighboring expert sets $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$, if $count_{\tau_{L}}(\Pi)<c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)=\tau_{L})<\Pr(E_{\tau_{L}}|\Pi)\cdot\delta^{\prime}.$ (40)
If $count_{\tau_{L}}(\Pi)\geq c_{min}$ then,
$\Pr(\mathcal{A}(\Pi)={\tau_{L}})\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})={\tau_{L}})$ (41)
Proof.
Let $\gamma^{(L)}=\top^{L}$.
$\Pr(\mathcal{A}(\Pi)=\tau_{L})=\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau}_{L})=\gamma^{(L)})\Pr(E_{\tau_{L}}|\Pi)$ (42)
If $count_{\tau_{L}}(\Pi)\geq c_{min}$, then using the fact that $\mathcal{A}_{2}$ is $(\varepsilon^{\prime},0)$-DP and Lemma 5.1, we can say that,
$\Pr(\mathcal{A}(\Pi)=\tau_{L})\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau_{L})$ (43)
Recall that the output of $\mathcal{A}_{2}$ is the output of a sequence of calls to $\mathcal{A}_{PQ}$ (Algorithm 1). The probability that $\mathcal{A}_{2}$ outputs the sequence $r$ can be upper-bounded by the probability of getting $\top$ on the $L^{th}$ call to $\mathcal{A}_{PQ}$. Formally,
$\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{L}})=\gamma^{(L)})\leq\Pr(\mathcal{A}_{PQ}(\Pi,\tau_{L},\hat{\theta},\varepsilon^{\prime})=\top)$ (44)
If $count_{\tau_{L}}(\Pi)<c_{min}$, then using $count_{\tau}(\Pi)<c_{min}<c_{min}/p_{min}=\theta$, (44) and Lemma 5.2, we get,
$\Pr(\mathcal{A}_{2}(\Pi,Q^{\tau_{L}})=\gamma^{(L)})\leq\Pr(\mathcal{A}_{PQ}(\Pi,\tau_{L},\hat{\theta},\varepsilon^{\prime})=\top)<\delta^{\prime}$ (45)
[…]
Combining 43 and 46, we get the statement of the lemma. ∎
A.4 Proof of Theorem 5.3
See 5.3
Proof.
Let $\mathcal{T}$ denote the set of all possible trajectory prefixes using $S,A$. From Lemma 5.4, we have, $\forall\tau\in\mathcal{T}$, and all pairs of neighboring sets of experts $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$,
$\Pr(\mathcal{A}(\Pi)=\tau)\leq e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau)+\Pr(E_{\tau}|\Pi)\cdot\delta^{\prime}$ (47)
where $\delta^{\prime}=\delta_{1}/(2LT)$.
To prove that $\mathcal{A}$ is $(2\varepsilon^{\prime},\delta_{1}/(2T))$-differentially private, we need to prove the following for any pair of neighbouring $\Pi,\Pi^{\prime}\subseteq\Pi_{p_{min}}$:
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}\leq\frac{\delta_{1}}{2T}$ (48)
Note that, for any fixed $k$,
$\sum_{\tau\in\{\tau:\tau\in\mathcal{T},|\tau|=k\}}\Pr(E_{\tau}|\Pi)=1$ (49)
where $|\tau|$ denotes the length of the trajectory prefix $\tau$.
$\sum_{\tau\in\mathcal{T}}\Pr(E_{\tau}|\Pi)=\sum_{k=1}^{L}\left[\sum_{\tau\in\{\tau:\tau\in\mathcal{T},|\tau|=k\}}\Pr(E_{\tau}|\Pi)\right]=L$ (50)
where $L$ is the length of a full trajectory. From equation 47, we have,
$\sum_{\tau\in\mathcal{T}}\max\{\Pr(\mathcal{A}(\Pi)=\tau)-e^{2\varepsilon^{\prime}}\Pr(\mathcal{A}(\Pi^{\prime})=\tau),0\}$ $\leq\sum_{\tau\in\mathcal{T}}\Pr(E_{\tau}|\Pi)\delta^{\prime}$ (51)
$=L\delta^{\prime}=\frac{\delta_{1}}{2T}$ (52)
This proves 48, and hence the statement of the theorem.
Appendix B EXPERIMENTAL DETAILS
B.1 Dataset Generation
To satisfy our assumptions, we enforce $p_{min}=0.02$ for all the environments.
Figure 5: Outline of Data Generation scheme; We train experts on multiple environments as described below.
Lunar Lander: Among the various variable which govern the dynamics of this environment, we chose to modify gravity (default = 9.8), wind power (default = 0.0) and turbulence (default = 0.0) to train different experts. We modified gravity over 10 equally spaced values from 9.0 to 11.0, wind power over 10 equally spaced values between 0.0 to 5.0, and similarly turbulence over 10 equally spaced values from 0.0 to 0.5, to generate 1000 environment settings. For each of these environment settings, we trained agents using the PPO algorithm for $61e$ 6 updates along with hyperparameter tuning of learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). For each environment setting, we chose the top 3 agents that performed the best on the corresponding environment, to form our 3000 experts.
CartPole: For CartPole, we chose to vary gravity, the magnitude of the force exerted when the cart is pushed and the mass of the cart. The default values of gravity, force magnitude and mass of the cart are 9.8, 10.0 and 1.0 respectively. We varied each control over 10 equally spaced values over a fixed range. For gravity, we fixed this range as (8.75, 11.0), for force magnitude as (9.0, 11.25) and for mass of the cart as (0.8, 1.25). Similar to above, for each of these 1000 environment settings, we trained agents using the PPO algorithm for $61e$ 6 updates while tuning the learning rate and batch size. Each agent is a neural network with hidden-layers of sizes (512, 256, 64). We chose the top 3 agents from each setting to form our 3000 experts.
Acrobot: For this environment, we varied the lengths of the 2 connected links and their masses. The default values of all these parameters is 1.0. We assign the same length to both the links in each setting. We vary this length, the mass of link 1 and the mass of link 2 over 10 equally spaced values across fixed ranges for each. We fix these ranges as (0.8, 1.2) for the length of the links, (0.9, 1.1) for the mass of link 1 and (0.9, 1.1) for the mass of link 2. The combination of these gives us 1000 environment settings. Again, we trained agents using PPO for $61e$ 6 updates over multiple learning rates and batch sizes and picked the top 3 agents from each setting to form our 3000 experts. Each agent is a neural network with hidden-layers of sizes (512, 256, 64).
HIV Treatment: In this domain, the 6-dimensional continuous state represents the concentrations of various cells and viruses. The actions correspond to the dosages of two drugs. The reward is a function of the health outcomes of the patient, and it also penalises the side-effects due to the quantity of drugs prescribed. Hence, the default reward function has 2 penalty terms, a penalty of 20000 on the dosage of drug 1, and a penalty of 2000 on the dosage of drug 2 (Ernst et al., (2006)). In the default setting of the environment, we assume that the treatment at each steps is applied for 20 days and that there are 50 time steps in an episode. To train experts, we vary the penalty on drug 1 over 50 equally spaced values from 18750 to 21250, and the penalty on drug 2 over 20 equally spaced values from 1850 to 2150. Another variation that we introduce is whether the efficacy of the administered drug is noisy or not. So we either add no noise to the drug efficacy, or we add gaussian noise with standard deviation 0.01. All these modifications give us 2000 different environment settings. We trained agents using PPO for $62e$ 6 updates on each of the environment settings and over multiple learning rates and batch sizes. Each agent is represented by a neural network with a single hidden layer of size 512. We then picked the top agents from each of the 2000 environment settings to create experts. We picked 1000 more top-performing agents from the 2nd best agents of all the environment settings to get 3000 experts in total.
Once we have chosen 3000 experts for any of the above environments, we modify each expert’s policy so that for each state, the action with the highest probability gets $(1-(|A|-1)\cdot p_{min})$ probability (where $|A|$ denotes the total number of actions), whereas all the other actions get $p_{min}$ probability of being executed. We then use these modified experts to sample 20 trajectories each from the default environment to form our offline dataset.
B.2 Expert-Level DP-SGD
We adapted DP-SGD to the expert-level privacy setting to form the baseline that we compare against. To achieve this, during gradient updates, we ensure that an expert contributes at most one transition to a batch. To construct a batch of size $b$, we randomly sample $b$ experts from the total $m$ experts, and then pick a transition from each selected expert. This is analogous to DP-SGD training on a dataset of size $m$. For privacy accounting, we use PLD accountant from the tensorflow_privacy library.
Algorithm 4 Expert-Level DP-SGD (outline)
1:Model Parameters $\phi$, Expert Datasets $D^{\Pi}$, Expert Policies $\Pi$, Loss function $\mathcal{L(\phi)}=\frac{1}{b}\sum_{i=1}^{b}\mathcal{L}(\phi,x_{i})$, Learning Rate $\eta$, Batch size $b$, Noise parameter $\sigma$, Clipping threshold $C$, Training steps $N$
2:Initialize $\phi_{0}$ randomly
3:for $t=1,\dots,N-1$ do
4:     Sample subset of experts $B$ from $\Pi$ with subsampling probability $\frac{b}{m}$
5:     For each $\pi_{i}\in B$ sample transition $x=(s,a,r,s^{\prime})$ from $D^{\pi_{i}}$ to form a batch of transitions $B_{trans}$
6:     For each $x_{i}\in B_{trans}$ compute gradients $g_{i}(\phi_{t})\leftarrow\nabla_{\phi_{t}}\mathcal{L}(\phi_{t},x_{i})$
7:     Scale gradients to fit clipping threshold $\bar{g}_{i}(\phi_{t})\leftarrow g_{i}(\phi_{t})/\text{max}\left(1,\frac{\left\lVert g_{i}(\phi_{t})\right\rVert_{2}}{C}\right)$
8:     Add gaussian noise to gradients $\tilde{g}_{t}\leftarrow\frac{1}{b}\left(\sum_{i=1}^{b}\bar{g}_{i}(\phi_{t})+\mathcal{N}(0,\sigma^{2}C^{2}\text{I})\right)$
9:     Update model parameters $\phi_{t+1}\leftarrow\phi_{t}-\eta\cdot\tilde{g}_{t}$
10:return $\phi_{T}$
Algorithm 4 describes a general outline for expert-level DP-SGD which we use as a baseline to compare our method against. The loss function here corresponds to the offline RL algorithm.
In our method, Algorithm 4 is incorporated in the training scheme in Algorithm 3 (lines 5 and 6) in the main paper. At each training step, with probability $p$, we execute lines 3-8 in Algorithm 4 to train on a batch of transitions. Otherwise, we sample a batch of transitions from $\mathcal{D}_{stable}^{\Pi}$ and perform a normal SGD update (this has no privacy cost). The privacy analysis for Algorithm 3 proceeds in a manner similar to that in standard DP-SGD (Abadi et al.,, 2016) with the exception of the subsampling probability which must be modified to be $q=p\cdot\frac{b}{m}$ (in the analysis only) to account for the randomness in picking the unstable dataset for sampling.
B.3 Hyperparameter Tuning
For all the environments, we assume that $p_{min}=0.02$. For LunarLander and CartPole, for $\varepsilon\leq 5$, we set $\varepsilon_{1}=0,\varepsilon_{2}=\varepsilon,\delta_{1}=0,\delta_{2}=\delta$ and resort to pure expert-level DP-SGD training since the privacy budget is not enough to get meaningful stable trajectories for training. Similarly for Acrobot, for $\varepsilon\leq 2.5$, we omit stable data collection and just do expert-level DP-SGD training. For HIV Treatment and for all other values of $\varepsilon$ for Lunar Lander and CartPole, we set $\varepsilon_{1}=\frac{3\varepsilon}{4},\varepsilon_{2}=\frac{\varepsilon}{4},\delta_{1}=\frac{9\delta}{10},\delta_{2}=\frac{\delta}{10}$. For Acrobot, we set $\varepsilon_{1}=\varepsilon,\varepsilon_{2}=0,\delta_{1}=\delta,\delta_{2}=0$.
We use Conservative Q-Learning (CQL) (Kumar et al., (2020)) for training on the offline data. For Lunar Lander, Acrobot and CartPole, we fix the batch size to 256, and the neural network hidden layer sizes to (256, 256). For HIV Treatment, we set the batch size as 64 and use a neural network of one hidden layer of size 512 to represent the Q-function. For DP-SGD updates, we clip all gradient norms to 1.0. We tune the following 3 hyperparameters for our method: learning rate $\eta$, probability of sampling from $\mathcal{D}_{unst}^{\Pi}$ during training ($p$), and DP-SGD noise multiplier which is the standard deviation of the gaussian noise applied to the gradients divided by the clipping threshold. We perform a grid search over all hyper-parameters. The search spaces for different hyperparameters are as follows: $\eta$ [0.0001, 0.001, 0.005, 0.01, 0.05], $p$ [0.9, 0.8, 0.5, 0.0], and noise multiplier - [10.0, 20.0, 30.0, 50.0, 60.0 80.0]. For each configuration of hyperparameters, we let the model train for as long as possible with the given value of $\varepsilon_{2}$. We report the average episodic return obtained over 10 evaluation runs spaced over the last 10000 steps of training.
Acrobot: We fix $p=0.0$, that is, we use only $\mathcal{D}_{stable}^{\Pi}$ for training. Over all values of $\varepsilon$, (5, 7.5, 10), $\eta$ = 0.005 gave the best performance. Since $p=0.0$, the noise multiplier is not required. For $\varepsilon=$ 10, 7.5, 5, we set $T$ (defined in Algorithm 2) equal to 25, 20, 20 respectively.
Lunar Lander: For $\varepsilon=10.0$, the best performing values of $\eta$, $p$ and noise multiplier are 0.005, 0.8, 60.0 respectively. For $\varepsilon=7.5$, the best performing hyperparameters, in the same order, are 0.005, 0.5, 80.0. For $\varepsilon=5.0,2.5$, we do pure expert-level DP-SGD training. For $\varepsilon=$ 10, 7.5, we set $T$ equal to 100, 62 respectively.
CartPole: For $\varepsilon=10.0$, the best performing values of $\eta$, $p$ and noise multiplier are 0.001, 0.9, 50.0 respectively. For $\varepsilon=7.5$, the best performing hyperparameters, in the same order, are 0.001, 0.8, 60.0. For lower values of $\varepsilon$, we do pure expert-level DP-SGD training. For $\varepsilon=$ 10, 7.5, we set $T$ equal to 25, 20 respectively.
HIV Treatment: We fix $p=0.0$, that is, we use only $\mathcal{D}_{stable}^{\Pi}$ for training (hence noise multiplier is not required). For $\varepsilon=10.0$, learning rate = 0.0001 gave the best performance. We use $T=50$.
For tuning the expert-level DP-SGD baseline, we vary $\eta$ and noise multiplier over the same ranges as specified above. We again let the model train for as long as possible with the given value of $\varepsilon$, and report the average episodic return obtained over 10 evaluation runs spaced over the last 10000 steps of training.